Using Charpit’s method, solve the equation:


zp² -y²p +y²q =0

1
Expert's answer
2022-01-25T07:03:15-0500

Using the Charpit's method, we shall solve PDE

"zp\u00b2-y\u00b2p+y\u00b2q"


Consider

"f(x,y,z,p,q)=0"


Given the PDE "zp\u00b2-y\u00b2p+y\u00b2q"

We have that

"f(x,y,z,p,q)" "=zp\u00b2-y\u00b2p+y\u00b2q=0"


We have the formula

"\displaystyle\frac{dp}{\frac{\partial f}{\partial x}+p\frac{\partial f}{\partial z}}=\frac{dq}{\frac{\partial f}{\partial y}+q\frac{\partial f}{\partial z}}=\frac{dz}{-p\frac{\partial f}{\partial p}-q\frac{\partial f}{\partial q}}=\frac{dx}{-\frac{\partial f}{\partial p}}=\frac{dy}{-\frac{\partial f}{\partial q}}"


It follows that

"\frac{\partial f}{\partial x}=0,\frac{\partial f}{\partial y}=-2yp+2yq,\frac{\partial f}{\partial z}=p²,\frac{\partial f}{\partial p}=2pz-2py,\frac{\partial f}{\partial q}=y²"


Substituting, we have

"\displaystyle \implies\frac{dp}{p³}=\frac{dq}{2py+2qy+p²q}=\frac{dz}{-p(2pz-y²)-q(y²)}=\frac{dy}{-y²}"



Consider

"\frac{dp}{p³}=\frac{dy}{-y²}"

We have that

"\displaystyle p^{-3}dp=-y^{-2}dy"


Integrating, we have

"\implies\int p^{-3}dp=\int -y^{-2}dy\\\implies\int p^{-3}dp=-\int y^{-2}dy"


"\implies \frac{p^{-2}}{-2}=y^{-1}+c"

Where "c" is the integrating constant.

"\implies \frac{1}{p²}=\frac{-2}{y}+c\\\implies y=c-2p²"

by making "p" the subject of the relation

"\implies p=\sqrt{\frac{c-y}{2}}"


Now, Substituting the value of "p" into the given PDE

We have

"\implies z(\frac{a-y}{2})-y²\sqrt{\frac{a-y}{2}}+y²q=0"


"\implies y²q=y²\sqrt{\frac{a-y}{2}}-z(\frac{a-y}{2})"


"\implies q=\sqrt{\frac{a-y}{2}}-\frac{z}{y²}(\frac{a-y}{2})"


we have been able to get values for "p\hspace{0.1cm}and\hspace{0.1cm}q"


NOW,

we know that "dz=p\hspace{0.05cm}dx+q\hspace{0.05cm}dy"

By substituting "p\hspace{0.1cm}and\hspace{0.1cm}q\hspace{0.1cm}in\hspace{0.1cm}dz"


We have

"dz=\sqrt{\frac{a-y}{2}}dx+\left\{\sqrt{\frac{a-y}{2}}-\frac{z}{y²}(\frac{a-y}{2})\right\}dy"


It follows that,


"\displaystyle dz=\int\sqrt{\frac{a-y}{2}}dx+\int \left\{\sqrt{\frac{a-y}{2}}-\frac{z}{y²}(\frac{a-y}{2})\right\}dy"


By integration, we have our final result to be


"\displaystyle z=\frac{x\sqrt{c-y}}{\sqrt{2}}+\left[\frac{\sqrt{2(c-y)}}{3}(y-c)+z\frac{c+y\hspace{0.05cm}log(y)}{2y}\right]"



Which is the solution of the PDE

"zp\u00b2-y\u00b2p+y\u00b2q"